OpenStudy (anonymous):

from a collection of 9 books, martin is asked to list his 5 favorites starting with his most favorite, second favorite, 3rd favorite and so on. How many ways can martin make his list?

OpenStudy (anonymous):

\[9\times 8\times 7\times 6\times 5\]

OpenStudy (kropot72):

\[nPr=\frac{n!}{(n-r)!}=\frac{9!}{4!}=9\times 8\times 7\times 6\times 5\]=15120 ways